OpenStudy (anonymous):

Find the x-intercept for -7x - 3y = -14

OpenStudy (bibby):

to find the x intercept, set y equal to 0 and solve

OpenStudy (tkhunny):

The Intercept Form is never far away. -7x - 3y = -14 7x + 3y = 14 \(\dfrac{x}{2} + \dfrac{y}{14/3} = 1\) And both s- and y-intercepts are staring at you.